A solid conducting sphere with an area of 16 m 2 has a total charge of 4 μC. Determine the electric field through the Gaussian surface.


1
Expert's answer
2022-02-09T06:29:12-0500

Given:

"A=\rm 16\:\rm m^2"

"q=\rm 4\:\mu C=4*10^{-6}\: C"

"\epsilon_0=8.85*10^{-12}\: \rm F/m"


The Gauss's law says

"\phi=\oint EdA=\frac{q}{\epsilon_0}"

For constant field

"EA=\frac{q}{\epsilon_0}"

"E=\frac{q}{\epsilon_0 A}=\frac{4*10^{-6}}{8.85*10^{-12}*16}=2.8*10^4\:\rm V/m"